--> A nurse needs to prepare 25 ml. of a solution that is 6.7% medicine (the rest is water). He or she has two bottles of solution: one that is 5% medicine, and one that is 10% medicine. How much from each bottle should he or she use?
here is what I tried:
0.5 = 5% = x = bottle 1
0.1 = 10% = y = bottle 2
6.7%/2 = 3.35%
0.0335(x) + 0.0335(y) = 0.067
0.0355(0.5) + 0.0335(0.1) = 0.067
0.01675 + 0.00335 = 0.201
I tried moving it to decimals but I had no luck.. Struggling out here in the cold....dark..equation..any suggestions?

Answer by lwsshak3(11628)

A nurse needs to prepare 25 ml. of a solution that is 6.7% medicine (the rest is water). He or she has two bottles of solution: one that is 5% medicine, and one that is 10% medicine. How much from each bottle should he or she use?
***
let x=amount of 5% medicine to use
25-x=amount of 10% medicine to use
5%x+10%(25-x)=6.7%*25
.05x+2.5-.10x=1.675
.05x=0.825
x=16.5
25-x=8.5
amount of 5% medicine to use=16.5 ml
amount of 10% medicine to use=8.5 ml